# Nested Skills in Labor Ecosystems: A Hidden Dimension of Human Capital

### v3

@@ Discussion
In increasingly complex, large teams, social skills become crucial when specialization requires workers to coordinate with team members possessing different specialized skills [Carlile2002, Wuchty2007, yoon2023, Neffkeeaax3370, Wu2019Science, Borner2018]. Our framework identifies and locates social skills embedded in the skill structure along with general and nested skills (SI Fig. LABEL:fig:social_skills-a), explaining their recent growth and significant role in wage premiums (see SI Fig. LABEL:fig:social_skills-b) [Liu2013, Deming2017, Borghans2014, Weinberger2014, Lindqvist2011, Kuhn2005, Deming2018, Kogan2021, VanderWouden2023, Evans2024]. Nevertheless, our results go beyond the contributions of social skills and managerial occupations to wage premiums, as the results are robust to their absence in analyses (see SI Sections LABEL:sec:social_skills and LABEL:sec:robustness_check:_no_managers). Therefore, social skills are valuable not just because of their role in sociality but because of their structural properties, serving as foundational building blocks of human capital to enable further valuable specialization and more complex organizations.
@@
For each skill, O*NET reports the required levels needed for workers of each occupation to perform their tasks. We call the distribution of the number of occupations that require skill at varying levels the level distribution. The shape of a skill’s level distribution illustrates its generality across occupations, shown in Fig. 1 (a). As such, we group skills by their similar distribution shapes by $k$-mean clustering algorithms with correlation metrics. Figure 1 (b) shows the characteristic shapes of each skill group. We provide three statistical tests for optimal $k$ and show the findings are qualitatively robust to some variations (see SI Sec. LABEL:supsec:skill_clustering). Throughout analyses, we mainly analyze the effects of general and specific skills to filter possible noises.
@@
The conditional probability that infers the directionality operates on binary values, but skill levels are recorded in continuous variables [0,7], which makes it hard to apply the conditional probability method. We use the disparity filter to extract a statistically significant presence/absence in an occupation-skill matrix [Serrano6483]. Parameters are chosen such that i) the rank of skill terms in the strength (from the weighted network) and degree (in the binary network) is preserved, ii) the rank of occupations’ skills of each category in the weighted network is preserved in the binary network. Supplementary Information Section LABEL:supsec:skill-occ discusses details and compares the state of data before and after the transformation.
We then calculate conditional probabilities of every pair of skills in the transformed (binary) matrix to infer dependence and directions between two skills, following [Jo2020]. We first account for the significance of conditional appearances, subject to a threshold, $z_{th}$. Here, $z_{th}$ is a threshold for the extent to which we eliminate chance from two skills being used in the same occupation. Given the significant skill pair conditional appearances, we estimate conditional probabilities $P(u|v)$ and $P(v|u)$. The direction of dependence $v\rightarrow u$ is set when $P(u|v)$ is substantially greater than $P(v|u)$, subject to a parameter $\alpha_{th}$, which is differentially weighted for each pair of skills so that it accounts for heterogeneous skill node degrees (see Eq. LABEL:eq:_a_th in SI Section LABEL:supsec:conditional_dependencies). The magnitude of the dependence is a parametric function of the difference between the conditional probabilities of observing $u$ and $v$, and the null model that accounts for the estimated number of shared occupations between them, given the degrees of $u$ and $v$.as shown by Eq. LABEL:eq:_dependency_weight in SI Section LABEL:supsec:conditional_dependencies. Figure 1 (d) broadly illustrates the intuition behind this methodology. Figure 2 (a) presents a backbone structure of the aggregated all skill pairs, where the edge weights follow the magnitudes of pairwise dependencies, as described above. Figure 2 (b) offers the full network. Please see SI Section LABEL:supsec:conditional_dependencies and [Jo2020] for the detailed procedures and choices of parameters and thresholds.
@@
Nestedness is a structural characteristic that describes interactions in an ecological system, where specialist species often interact with a subset of generalists. Unlike ecological systems, however, SI-Fig. LABEL:fig:occ_skill_nestedness_mat shows the skill-occupation matrix is a noisy nested structure far from the perfect upper-left triangle when sorted by marginal totals (fills). This imperfect nested structure may account for the constraints on occupations (limited carrying capacity), introducing severe competition between skill species. Indeed, SI-Fig. LABEL:fig:occ_vs_skill_importance_avg_cos shows, unlike broad skill generality, the occupation’s scope is narrowly distributed, indicating that the total amount of skill levels embodied in an occupation is not much different from each other, regardless of how much they are paid and how advanced education is needed (see SI Sec. LABEL:suppsec:nestedness).
@@
Therefore, we quantify a skill’s contribution to the nested structure, i.e., nested score, $c_{s}$, defined as a deviation from a null model where the edges of a focal node $s$ to occupations are randomly reassigned, that is, $c_{s}=(N-<N^{\ast}_{s}>)/{\sigma_{N^{\ast}_{s}}}$. $N$ is a nestedness score, and $<N^{\ast}_{s}>$ and $\sigma_{N^{\ast}_{s}}$ are the means and standard deviation derived from the null model [Saavedra2011]. For each focal skill $s$, we run 5,000 iterations [SergeiMaslov2002]. We employ the overlap index, checkerboard score, Temperature, and NODF, nestedness scores commonly used in ecology, to quantify nestedness $N$ [write1992, Stone1990, Atmar1993, Almeida-neto2008]. In addition, we only consider skill’s contribution and do not occupation’s contribution. To obtain discrete categorizes, any non-general skill with $c_{s}>0$ is called ”nested” skills, and ”un-nested” otherwise. The resulting skill categories are shown in Fig. 3. The detailed allocation of skills to these categories are outlined in SI Table LABEL:tab:skill_split_C_nestedness, and SI Sec. LABEL:suppsec:nestedness offers details and robustness checks.
@@
Demographic analysis: CPS microdata also include gender and race/ethnicity demographic information. We chose four categories, Whites, Blacks, Asians, and Hispanic, as they are the bulk of the sample, and any individuals of Hispanic background are included in that category for Fig. 6. To avoid attrition and early retirement, we include only full-time workers employed at the time of the survey, earning at least \$10,000 annually, and between 18 and 55. For each demographic category, the average skill level is calculated for their occupational composition. The microdata records individuals’ wages and the number of hours worked. We adjust wages for inflation and account for the number of hours worked to compute an adjusted weekly wage, which is readily comparable across the population. The race/ethnic disparities in Fig. 6 are a ratio of each demographic quantity (general level, nested level, unnested levels, education, and weekly wages) to those of White workers, following [Tong2021] identifying a dominant social group, a social group if it is at least 1.5 times more likely to be employed in the focal occupation. Likewise, the gender gap within each race/ethnicity is measured as a ratio of those quantities to those of male workers. Because we do not have a matched sample, we obtain 95% confidence intervals by random sub-sampling. In each iteration, we take 10% of the subpopulation of interest, for instance, Asian male and Asian female workers, and estimate all corresponding measures. Repeating this sampling and estimation process in 10,000 iterations, we obtain the distribution for each estimation and derive the 95% confidential intervals. The skill, education, and wage estimations of Fig. 6 average over the years. Supplementary Figs. LABEL:fig:Temporal_Race_Gaps_-_Skills,_Education,_Wages and LABEL:fig:Temporal_Gender_Gaps_-_Skills,_Education,_Wages capture temporal patterns of these factors, exhibiting the gaps have narrowed over time. In addition, SI Figs. LABEL:fig:Skill_Age_Gender_Race_Trends and LABEL:fig:Skill_Age_Gender_Race_Trends_-_year_effects show the skill differentials between male and female workers that start around the age of 30 (main Fig. 4), manifest across racial and ethnic groups.
@@
We exclude job transitions shorter than one year or within an occupation (i.e., moving from one company to another without changing the occupation) for our primary analyses. The decision to remove such occupations arises from the oddity we observed in most such jobs. For instance, various janitors or models became CEOs immediately or with overlapping periods. Nevertheless, our findings are robust to this decision (see SI Sec LABEL:supsec:_skill_dependencies_and_age for details).

### v4

@@ Discussion
In increasingly complex, large teams, social skills become crucial when specialization requires workers to coordinate with team members possessing different specialized skills [Carlile2002, Wuchty2007, yoon2023, Neffkeeaax3370, Wu2019Science, Borner2018]. Our framework identifies and locates social skills embedded in the skill structure along with general and nested skills (SI Fig. S63-a), explaining their recent growth and significant role in wage premiums (see SI Fig. S63-b) [Liu2013, Deming2017, Borghans2014, Weinberger2014, Lindqvist2011, Kuhn2005, Deming2018, Kogan2021, VanderWouden2023, Evans2024]. Nevertheless, our results go beyond the contributions of social skills and managerial occupations to wage premiums, as the results are robust to their absence in analyses (see SI Sections 9.2 and 9.1). Therefore, social skills are valuable not just because of their role in sociality but because of their structural properties, serving as foundational building blocks of human capital to enable further valuable specialization and more complex organizations.
@@
For each skill, O*NET reports the required levels needed for workers of each occupation to perform their tasks. We call the distribution of the number of occupations that require skill at varying levels the level distribution. The shape of a skill’s level distribution illustrates its generality across occupations, shown in Fig. 1 (a). As such, we group skills by their similar distribution shapes by $k$-mean clustering algorithms with correlation metrics. Figure 1 (b) shows the characteristic shapes of each skill group. We provide three statistical tests for optimal $k$ and show the findings are qualitatively robust to some variations (see SI Sec. 1). Throughout analyses, we mainly analyze the effects of general and specific skills to filter possible noises.
@@
The conditional probability that infers the directionality operates on binary values, but skill levels are recorded in continuous variables [0,7], which makes it hard to apply the conditional probability method. We use the disparity filter to extract a statistically significant presence/absence in an occupation-skill matrix [Serrano6483]. Parameters are chosen such that i) the rank of skill terms in the strength (from the weighted network) and degree (in the binary network) is preserved, ii) the rank of occupations’ skills of each category in the weighted network is preserved in the binary network. Supplementary Information Section 3.1 discusses details and compares the state of data before and after the transformation.
We then calculate conditional probabilities of every pair of skills in the transformed (binary) matrix to infer dependence and directions between two skills, following [Jo2020]. We first account for the significance of conditional appearances, subject to a threshold, $z_{th}$. Here, $z_{th}$ is a threshold for the extent to which we eliminate chance from two skills being used in the same occupation. Given the significant skill pair conditional appearances, we estimate conditional probabilities $P(u|v)$ and $P(v|u)$. The direction of dependence $v\rightarrow u$ is set when $P(u|v)$ is substantially greater than $P(v|u)$, subject to a parameter $\alpha_{th}$, which is differentially weighted for each pair of skills so that it accounts for heterogeneous skill node degrees (see Eq. 6 in SI Section 3). The magnitude of the dependence is a parametric function of the difference between the conditional probabilities of observing $u$ and $v$, and the null model that accounts for the estimated number of shared occupations between them, given the degrees of $u$ and $v$.as shown by Eq. 7 in SI Section 3. Figure 1 (d) broadly illustrates the intuition behind this methodology. Figure 2 (a) presents a backbone structure of the aggregated all skill pairs, where the edge weights follow the magnitudes of pairwise dependencies, as described above. Figure 2 (b) offers the full network. Please see SI Section 3 and [Jo2020] for the detailed procedures and choices of parameters and thresholds.
@@
Nestedness is a structural characteristic that describes interactions in an ecological system, where specialist species often interact with a subset of generalists. Unlike ecological systems, however, SI-Fig. S6 shows the skill-occupation matrix is a noisy nested structure far from the perfect upper-left triangle when sorted by marginal totals (fills). This imperfect nested structure may account for the constraints on occupations (limited carrying capacity), introducing severe competition between skill species. Indeed, SI-Fig. S5 shows, unlike broad skill generality, the occupation’s scope is narrowly distributed, indicating that the total amount of skill levels embodied in an occupation is not much different from each other, regardless of how much they are paid and how advanced education is needed (see SI Sec. 2).
@@
Therefore, we quantify a skill’s contribution to the nested structure, i.e., nested score, $c_{s}$, defined as a deviation from a null model where the edges of a focal node $s$ to occupations are randomly reassigned, that is, $c_{s}=(N-<N^{\ast}_{s}>)/{\sigma_{N^{\ast}_{s}}}$. $N$ is a nestedness score, and $<N^{\ast}_{s}>$ and $\sigma_{N^{\ast}_{s}}$ are the means and standard deviation derived from the null model [Saavedra2011]. For each focal skill $s$, we run 5,000 iterations [SergeiMaslov2002]. We employ the overlap index, checkerboard score, Temperature, and NODF, nestedness scores commonly used in ecology, to quantify nestedness $N$ [write1992, Stone1990, Atmar1993, Almeida-neto2008]. In addition, we only consider skill’s contribution and do not occupation’s contribution. To obtain discrete categorizes, any non-general skill with $c_{s}>0$ is called ”nested” skills, and ”un-nested” otherwise. The resulting skill categories are shown in Fig. 3. The detailed allocation of skills to these categories are outlined in SI Table LABEL:tab:skill_split_C_nestedness, and SI Sec. 2 offers details and robustness checks.
@@
Demographic analysis: CPS microdata also include gender and race/ethnicity demographic information. We chose four categories, Whites, Blacks, Asians, and Hispanic, as they are the bulk of the sample, and any individuals of Hispanic background are included in that category for Fig. 6. To avoid attrition and early retirement, we include only full-time workers employed at the time of the survey, earning at least \$10,000 annually, and between 18 and 55. For each demographic category, the average skill level is calculated for their occupational composition. The microdata records individuals’ wages and the number of hours worked. We adjust wages for inflation and account for the number of hours worked to compute an adjusted weekly wage, which is readily comparable across the population. The race/ethnic disparities in Fig. 6 are a ratio of each demographic quantity (general level, nested level, unnested levels, education, and weekly wages) to those of White workers, following [Tong2021] identifying a dominant social group, a social group if it is at least 1.5 times more likely to be employed in the focal occupation. Likewise, the gender gap within each race/ethnicity is measured as a ratio of those quantities to those of male workers. Because we do not have a matched sample, we obtain 95% confidence intervals by random sub-sampling. In each iteration, we take 10% of the subpopulation of interest, for instance, Asian male and Asian female workers, and estimate all corresponding measures. Repeating this sampling and estimation process in 10,000 iterations, we obtain the distribution for each estimation and derive the 95% confidential intervals. The skill, education, and wage estimations of Fig. 6 average over the years. Supplementary Figs. S50 and S51 capture temporal patterns of these factors, exhibiting the gaps have narrowed over time. In addition, SI Figs. S52 and S53 show the skill differentials between male and female workers that start around the age of 30 (main Fig. 4), manifest across racial and ethnic groups.
@@
We exclude job transitions shorter than one year or within an occupation (i.e., moving from one company to another without changing the occupation) for our primary analyses. The decision to remove such occupations arises from the oddity we observed in most such jobs. For instance, various janitors or models became CEOs immediately or with overlapping periods. Nevertheless, our findings are robust to this decision (see SI Sec 4 for details).